# MagicTailor: Component-Controllable Personalization in Text-to-Image Diffusion Models

### v2

@@ A.6 Compared Methods
In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) [11], DreamBooth-LoRA (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38]. We select these methods because TI, DB, and CD are three representatives of personalization frameworks and BAS and CLiC are highly relevant to learning fine-grained elements from reference images. For TI, DB, and CD, we use the third-party implementation in Diffusers [8]. For BAS, we use the official implementation. For CLiC, we reproduce it following the resource paper as the official code is not released. Unless otherwise specified, method-specific configurations are set up by following their resource papers or Diffusers. We empirically adjust the learning rate of CD and CLiC to 1e-4 and 5e-5 respectively, because they perform very poorly with the original learning rates. For a fair and meaningful comparison, these methods should be adapted to our task setting with minimal modification. Therefore, for those methods adopting a vanilla diffusion loss, we integrate the masked diffusion loss into them while using the same segmentation masks from MagicTailor.
Figure 10: Compatibility with different backbones. We equip MagicTailor with SD 1.5 [35], SD 2.1 [35], and SDXL [27]. The results show that MagicTailor can be generalized to multiple backbones, and a better backbone could provide better performance.
@@ B.1 Detailed Text-Guided Generation
One might wonder if component-controllable personalization can be accomplished by providing detailed textual descriptions to the T2I model. To investigate this, we separately feed the reference images of the concept and component into GPT-4o [17] to obtain detailed textual descriptions for them. The text prompt we used is “Please detailedly describe the <concept/component> of the upload images in a parapraph”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 [35] to generate the corresponding images. Some examples for a qualitative comparison are shown in Fig. 9. As we can see, such an approach cannot achieve satisfactory results, because it is hard to guarantee that visual semantics can be completely expressed by using the combination of text tokens. In contrast, our MagicTailor is able to accurately learn the desired visual semantics of the concept and component from reference images, and thus lead to consistent and excellent generation in this tough task.
@@ B.2 Commercial Models
It is also worth exploring whether existing commercial models, which can understand and somehow generate both text and images by themselves or other integrated tools, are capable of handling component-controllable personalization. We choose two widely recognized commercial models, GPT-4o [17] and Gemini 1.5 Flash [43], for a qualitative comparison. First, we separately feed the reference images of the concept and component into them, along with the text prompt of “The uploaded images contain a special instance of the <concept/component>, please mark it as #<concept/component>”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we instruct them to perform image generation, using the text prompt of “Please generate images containing #<concept> with #<component>”, where “<concept>” and “<component>” are replaced with the category label of the concept and component, respectively. As shown in Fig. 12, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component. Whereas, our MagicTailor achieves superior results in component-controllable personalization, using a dedicated framework designed for this task. It demonstrates that, even though large commercial models are able to tackle multiple general tasks, there is also plenty of room for the community to explore specialized tasks for real-world applications.
Figure 11: Effectiveness on difficult training pairs. We show the results of two hard cases involving large geometric discrepancy and cross-domain interactions, showing that MagicTailor can effectively handle such challenging scenarios.
Appendix C Generalizability Analysis
As discussed in Sec. A, we follow a strict and standardized setup to conduct performance evaluation for all the methods in our experiments. In this section, we perform a comprehensive generalizability analysis for MagicTailor by extending the setup to more flexible scenarios. Specifically, the analysis focuses on the model perspective, where we explore MagicTailor’s compatibility with different backbones (Sec. C.1), and the data perspective, where we study its effectiveness on difficult training pairs (Sec. C.2) and complex text prompts (Sec. C.3).
@@ Appendix C Generalizability Analysis
Figure 13: Effectiveness on complex text prompts. We present qualitative results generated with complex text prompts. In addition to those well-categorized text prompts, our MagicTailor can also follow more complex ones to generate text-aligned images.
C.1 Compatibility with Different Backbones
We use Stable Diffusion 2.1 (SD 2.1) [35], which is widely adopted by the compared methods, as our backbone in the main paper. It is worth noting that our MagicTailor is a general method also suitable for other T2I diffusion models, because only input images and the denoising paradigm are reformulated. To evaluate the compatibility, we employ MagicTailor in other backbones, including Stable Diffusion 1.5 (SD 1.5) [35] and Stable Diffusion XL (SDXL) [27]. As shown in Fig. 10, our MagicTailor can also achieve satisfactory results when cooperating with various backbones. Notably, we only change the LoRA rank for adapting to different learning capabilities, and directly use the original hyperparameter values, which shows the remarkable generalizability of MagicTailor.
C.2 Effectiveness on Difficult Training Pairs
Appropriate processing for reference images, such as filtering, scaling, and cropping, can help to construct a better training pair for personalization. Unfortunately, not all training pairs are ideal in real-world scenarios. Therefore, here we investigate the performance of MagicTailor when handling difficult training pairs. Specifically, we consider two hard cases. The first is a pair with large geometric discrepancy, such as “<person>” in an upper body portrait and “<hair>” in a profile photo. The second is a pair with cross-domain interactions, such as “<person>” and “<ear>” of dogs. As shown in Fig. 11, even facing these hard cases, MagicTailor is also capable of personalizing the target concept and component with high fidelity.
C.3 Effectiveness on Complex Text Prompts
In our experiments, we use well-categorized text prompts in four aspects, including recontextualization, restylization, interaction, and property modification, for qualitative and quantitative comparisons. Here we further evaluate MagicTailor’s performance on other complex text prompts involving more complicated contexts and elements. As shown in Fig. 13, our MagicTailor can also effectively generate text-aligned images when performing fidelity personalization, showing its ability to handle diverse user needs.
@@ C.3 Effectiveness on Complex Text Prompts
Figure 14: Ablation of DM-Deg via replacement with SID. We compare our DM-Deg with SID [18] that aims to produce informative prompts for training. Besides, we also present baseline (i.e., removing DM-Deg from MagicTailor) results for reference. This comparison indicates the effectiveness of our DM-Deg in addressing semantic pollution.
Appendix D Additional Ablation Studies
D.1 Effectiveness of DM-Deg over Using Informative Training Prompts
One might be curious about whether it is not necessary to employ the proposed DM-Deg, but perhaps to use informative text prompts during training to provide textual prior knowledge for learning the target concept and component. To investigate this, we use Selectively Informative Description (SID) [18] with GPT-4o [17] to construct text prompts for the target concept and component, and then use them for training. As shown in Fig. 14, such an approach cannot address semantic pollution well, where unwanted visual semantics still affect the personalized concept. In contrast, DM-Deg effectively prevents semantic pollution by dynamically perturbing those undesired visual semantics, verifying its remarkable significance in this task.
D.2 Necessity of Warm-Up Training
@@
In MagicTailor, we start with a warm-up phase for the T2I model to preliminarily inject the knowledge for the subsequent phase of DS-Bal. Here we investigate the necessity of such a warm-up phase for generation performance. In Tab. 7, when removing the warm-up phase, even though MagicTailor could obtain slight improvement in text alignment, it severely suffers from the huge drop in identity fidelity. This is because such a scheme makes it difficult to construct a decent momentum denoising U-Net for DS-Bal. Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance, which is attributed to the knowledge reserved from warm-up.
@@
Figure 15: Ablation of the number of reference images. We present qualitative results to show that MagicTailor can still achieve satisfactory performance when provided only 1 or 2 reference image(s) per concept and component.
D.3 Performance on Different Numbers of Reference Images
As described in Appendix A, each concept/component contains 3 reference images in a concept-component pair of the dataset. Here we change the number of reference images to analyze the performance variation of MagicTailor. The qualitative results are presented in Fig. 15. When the number of reference images is reduced, MagicTailor can still show satisfactory performance. This also demonstrates that, while more reference images could lead to better generalization ability, one reference image per concept/component is enough to obtain a decent result with our MagicTailor.
D.4 Robustness on Linking Words
@@
Figure 16: Ablation of linking words. We present qualitative results generated with different linking words, demonstrating the robustness of MagicTailor.
Appendix E More Qualitative Results
In Fig. 17 & Fig. 18, we provide more evaluation images for a substantial qualitative comparison. It can be clearly observed that semantic pollution remains an intractable problem for these compared methods. This is due to the leak of an effective mechanism to alleviate the T2I model’s perception for these semantics. To address this, our MagicTailor utilizes DM-Deg to dynamically perturb undesired visual semantics during the learning phase, and thus achieve better performance. On the other hand, the compared methods are also severely influenced by semantic imbalance, resulting in overemphasis or even overfitting on the concept or component. This is because the inherent imbalance of visual semantics complicates the learning process. In response to this issue, our MagicTailor applies DS-Bal to balance the learning of visual semantics, effectively showcasing its prowess in this tough task.

### v3

@@ A.6 Compared Methods
In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) [Gal et al., 2022], DreamBooth-LoRA (DB) [Ruiz et al., 2023], Custom Diffusion (CD) [Kumari et al., 2023], Break-A-Scene (BAS) [Avrahami et al., 2023], and CLiC [Safaee et al., 2024]. We select these methods because TI, DB, and CD are three representatives of personalization frameworks and BAS and CLiC are highly relevant to learning fine-grained elements from reference images. For TI, DB, and CD, we use the third-party implementation in Diffusers [Face, 2022]. For BAS, we use the official implementation. For CLiC, we reproduce it following the resource paper as the official code is not released. Unless otherwise specified, method-specific configurations are set up by following their resource papers or Diffusers. We empirically adjust the learning rate of CD and CLiC to 1e-4 and 5e-5 respectively, because they perform very poorly with the original learning rates. For a fair and meaningful comparison, these methods should be adapted to our task setting with minimal modification. Therefore, for those methods adopting a vanilla diffusion loss, we integrate the masked diffusion loss into them while using the same segmentation masks from MagicTailor.
@@ B.1 Detailed Text-Guided Generation
One might wonder if component-controllable personalization can be accomplished by providing detailed textual descriptions to the T2I model. To investigate this, we separately feed the reference images of the concept and component into GPT-4o [Hurst et al., 2024] to obtain detailed textual descriptions for them. The text prompt we used is “Please detailedly describe the <concept/component> of the upload images in a parapraph”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 [Rombach et al., 2022] to generate the corresponding images. Some examples for a qualitative comparison are shown in Fig. 13. As we can see, such an approach cannot achieve satisfactory results, because it is hard to guarantee that visual semantics can be completely expressed by using the combination of text tokens. In contrast, our MagicTailor is able to accurately learn the desired visual semantics of the concept and component from reference images, and thus lead to consistent and excellent generation in this tough task.
@@ B.2 Commercial Models
It is also worth exploring whether existing commercial models, which can understand and somehow generate both text and images by themselves or other integrated tools, are capable of handling component-controllable personalization. We choose two widely recognized commercial models, GPT-4o [Hurst et al., 2024] and Gemini 1.5 Flash [Team et al., 2023], for a qualitative comparison. First, we separately feed the reference images of the concept and component into them, along with the text prompt of “The uploaded images contain a special instance of the <concept/component>, please mark it as #<concept/component>”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we instruct them to perform image generation, using the text prompt of “Please generate images containing #<concept> with #<component>”, where “<concept>” and “<component>” are replaced with the category label of the concept and component, respectively. As shown in Fig. 14, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component. Whereas, our MagicTailor achieves superior results in component-controllable personalization, using a dedicated framework designed for this task. It demonstrates that, even though large commercial models are able to tackle multiple general tasks, there is also plenty of room for the community to explore specialized tasks for real-world applications.
@@
Figure 15: Ablation of DM-Deg via replacement with SID. We compare our DM-Deg with SID [Kim et al., 2024] that aims to produce informative prompts for training. Besides, we also present baseline (i.e., removing DM-Deg from MagicTailor) results for reference. This comparison indicates the effectiveness of our DM-Deg in addressing semantic pollution.
Appendix C More Ablation Studies and Analysis
C.1 Dynamic Intensity Matters
C.2 Momentum Denoising U-Net as a Good Regularizer
C.3 Necessity of Warm-Up Training
@@
In MagicTailor, we start with a warm-up phase for the T2I model to preliminarily inject the knowledge for the subsequent phase of DS-Bal. Here we investigate the necessity of such a warm-up phase for generation performance. In Tab. 7, when removing the warm-up phase, even though MagicTailor could obtain slight improvement in text alignment, it severely suffers from the huge drop in identity fidelity. This is because such a scheme makes it difficult to construct a decent momentum denoising U-Net for DS-Bal. Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance due to the knowledge reserved from warm-up.
Mask-Out Startegy 0.270 0.818 0.760 0.375
Fixed ($\alpha=0.4$) 0.270 0.849 0.800 0.297
Fixed ($\alpha=0.6$) 0.271 0.845 0.794 0.310
Fixed ($\alpha=0.8$) 0.271 0.846 0.796 0.305
Linear (Ascent) 0.270 0.846 0.797 0.307
Linear (Descent) 0.261 0.851 0.802 0.300
Dynamic ($\gamma=8$) 0.266 0.850 0.806 0.289
Dynamic ($\gamma=16$) 0.268 0.854 0.813 0.282
Dynamic ($\gamma=64$) 0.271 0.852 0.812 0.283
Dynamic (Ours) 0.270 0.854 0.813 0.279
C.4 Effectiveness of DM-Deg over Using Informative Training Prompts
One might be curious about whether it is not necessary to employ the proposed DM-Deg, but perhaps to use informative text prompts during training to provide textual prior knowledge for learning the target concept and component. To investigate this, we use Selectively Informative Description (SID) [Kim et al., 2024] with GPT-4o [Hurst et al., 2024] to construct text prompts for the target concept and component, and then use them for training. As shown in Fig. 15, such an approach cannot address semantic pollution well, where unwanted visual semantics still affect the personalized concept. In contrast, DM-Deg effectively prevents semantic pollution by dynamically perturbing those undesired visual semantics, verifying its remarkable significance in this task.
Fixed ($\beta=0$) 0.268 0.850 0.803 0.293
Fixed ($\beta=1$) 0.270 0.851 0.808 0.286
Momentum ($\beta=0.5$) 0.268 0.850 0.805 0.290
Momentum ($\beta=0.9$) 0.269 0.850 0.808 0.288
Momentum (Ours) 0.270 0.854 0.813 0.279
@@ C.4 Effectiveness of DM-Deg over Using Informative Training Prompts
C.5 Robustness on Linking Words
@@
Figure 16: Ablation of linking words. We present qualitative results generated with different linking words in text prompts, demonstrating the robustness of MagicTailor.
Appendix D More Qualitative Results
In Fig. 17 & Fig. 18, we provide more evaluation images for a substantial qualitative comparison. It can be clearly observed that semantic pollution remains an intractable problem for these compared methods. This is due to the leak of an effective mechanism to alleviate the T2I model’s perception for these semantics. To address this, our MagicTailor utilizes DM-Deg to dynamically perturb undesired visual semantics during the learning phase, and thus achieve better performance. On the other hand, the compared methods are also severely influenced by semantic imbalance, resulting in overemphasis or even overfitting on the concept or component. This is because the inherent imbalance of visual semantics complicates the learning process. In response to this issue, our MagicTailor applies DS-Bal to balance the learning of visual semantics, effectively showcasing its prowess in this tough task. In summary, the proposed MagicTailor effectively addresses both semantic pollution and semantic imbalance through its innovative techniques, DM-Deg and DS-Bal, respectively. These advancements demonstrate its superiority in handling complex visual semantics and achieving remarkable performance in this challenging task.